OpenStudy (anonymous):

what are the x- and y-intercepts of y= (x+2)√x show work?

OpenStudy (zale101):

Solving for x intercepts \(\bf you~set~y=0\) in the equation Solving for y intercepts \(\bf you~set~x=0\) in the equation

OpenStudy (anonymous):

what would that give you though

OpenStudy (zale101):

Solving for y intercepts \(\bf you~set~x=0\) in the equation that means \(\bf y= ((0)+2)\sqrt{(0)}\) what would you get for y?

OpenStudy (anonymous):

you would get 0

OpenStudy (zale101):

Yes, that means the y-intercept is?

OpenStudy (anonymous):

i know that for the answer it would be (0,0) and (-2,0) for the x intercepts and (0,0) for the y i just need to know how to do it

OpenStudy (zale101):

Y intercept is (0,0) Solving for x intercepts \(\bf you~set~y=0\) that means \(\bf 0=(x+2)(\sqrt{x})\) \(\color{red}{(x+2)}=0\) and \(\color{blue}{(\sqrt{x})}=0\) \(\color{red}{x=-2}\) and \(\color{blue}{x=0}\) so the x intercept will lie on (-2,0) and (0,0)

OpenStudy (zale101):

what part that you need more clarification of?

OpenStudy (zale101):

i'll be glad to clarify

OpenStudy (anonymous):

yes that makes sense. could you help me with another one?

OpenStudy (zale101):

Sure

OpenStudy (anonymous):

how do you show that there is x-axis symmetry for x=√(9-y^2)

OpenStudy (anonymous):

x=sqrt(9-y^2)

OpenStudy (anonymous):

@Zale101

OpenStudy (zale101):

to test for symmetry for f(x), you change the input positive x to negative x so f(-x).

OpenStudy (anonymous):

how would you do that

OpenStudy (anonymous):

@Zale101

OpenStudy (zale101):

\(f(y)=\sqrt{(9-y^2)}\) \(f(-y)=\sqrt{(9-(-y)^2)}\) y is square rooted, that indicates that y is positive no matter what number we put there. That means, \(f(-y)\) has the same function of \(f(y)\). If i were to reflect the graph about the y axis, it would still look the same as the original equation. The name of this symmetry is y axis symmetry. For more proof, let's see the graph of the equation. 

OpenStudy (zale101):